Activity 2



Find the rejection region for each hypothesis test based on the information given.



1 H₂ μ = 121



Hah > 121 Η μ=986



2. Ho:μ=986 Ị Hai =27



HaiH <27



a= 0.01



a= 0.05 a= 0.05



n=34



known unknown



n=25 n=12



a known

1
Expert's answer
2022-05-27T14:32:42-0400

1.

"H_0: \mu=121"

"H_a:\mu>121"

"\sigma" known, "n=34, \alpha=0.01"

This corresponds to a right-tailed test, for which a z-test for one mean, with known population standard deviation will be used.


"z\ge2.3263"

2.

"H_0: \mu=986"

"H_a:\mu\not=986"

"\sigma" unknown, "n=25, \alpha=0.5"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

"df=n-1=24" degrees of freedom


"|t|\ge2.0639"

3.

"H_0: \mu=27"

"H_a:\mu<27"

"\sigma" known, "n=12, \alpha=0.05"

This corresponds to a left-tailed test, for which a z-test for one mean, with known population standard deviation will be used.


"z\u2264-1.6449"